OpenStudy (anonymous):

(3-4 (-2+h)^2) - (3-4 (-2)^2)/h

OpenStudy (anonymous):

\[(3-4 (-2+h)^2)- \frac{ (3-4 (-2)^2) }{ h }\]

OpenStudy (anonymous):

what do you think is the result?

OpenStudy (anonymous):

@thatsme

OpenStudy (anonymous):

Well, the answer is 16h-4h^2/h. But, when I work it out I seem to be missing a step because I end up with a different answer.

OpenStudy (anonymous):

\[-4 (h-2)^2+\frac{ 13 }{ h }+3\] well this is the answer I think....

OpenStudy (anonymous):

are there any choices?

OpenStudy (anonymous):

No, the answer I provided is what the homework answer key says.

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

hold on let me give you some equations.....It's gonna take a little while...

OpenStudy (anonymous):

\[\frac{ -(4 h^3-16 h^2+13 h-13) }{ h }\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

h = 3.3154

OpenStudy (anonymous):

\[h = 0.34230-0.92903 i\]

OpenStudy (anonymous):

this one is gonna take a little long while....

OpenStudy (anonymous):

\[\frac{ d }{ dh} | (3-4 (-2+h)^2)-\frac{ (3-4 (-2)^2) }{ h}| = -\frac{ 13 }{ h^2 }-8 h+16\]

OpenStudy (anonymous):

\[3-4 (-2+h)^2+\frac{ 13 }{ h } =\]

OpenStudy (anonymous):

*sweats and breathes heavily*

OpenStudy (anonymous):

\[\sum_{n= - \infty}^{\infty}\]

OpenStudy (anonymous):

THERE IS THE EQUATIOn!!!

hero (hero):

@thatsme, can you write out what the problem looks like using the draw button?